OpenStudy (barrelracer17):

Find the coordinates for the midpoint of the segment with endpoints given. (3, 5) and (-2, 0)

pooja195 (pooja195):

\[\huge~\rm~m=\frac{ x_1+x_2 }{ 2 },\frac{ y_1+y_2 }{ 2 }\]

pooja195 (pooja195):

Use that and plug in the numbers :)

pooja195 (pooja195):

\[\huge~\rm~m=\frac{ 3-2 }{ 2 },\frac{ 5+0 }{ 2 }\]

pooja195 (pooja195):

Im sure you can take it from there :)

OpenStudy (barrelracer17):

Thank. You so much

pooja195 (pooja195):

No problem :) but you still need to find x and y

OpenStudy (barrelracer17):

Do I use that to find x and y?

pooja195 (pooja195):

\[\huge~\rm~m=\frac{ 3-2 }{ 2 },\frac{ 5+0 }{ 2 }\] \(\huge\color{red}{(x,y)}\)

pooja195 (pooja195):

whats 3-2?

OpenStudy (barrelracer17):

1

pooja195 (pooja195):

\(\huge\color{red}{x=1/2}\)

pooja195 (pooja195):

5+0= ?

OpenStudy (barrelracer17):

5

pooja195 (pooja195):

\(\huge\color{red}{y=5/2}\)

pooja195 (pooja195):

whats the answer?

OpenStudy (barrelracer17):

Well 1/2 is .5 and 5/2 is 2.5 according to my calculator

pooja195 (pooja195):

keep it as a fraction

OpenStudy (barrelracer17):

Thank you so much

pooja195 (pooja195):

:) Welcome to OpenStudy :)